OpenStudy (anonymous):

Solve the given system, using the substitution method. y=3x–4 9x–3y=14 A. There is no solution. B. (12, 14) C. (14, 12) D. There are an infinite number of solutions.

OpenStudy (anonymous):

So do you know what the first step is?

OpenStudy (anonymous):

Its A. By the way

OpenStudy (anonymous):

y=3x-4 9x-3y=14 sub in 3x-4 in for y in the 2nd equation 9x-3(3x-4)=14 9x-9x+12=14 ----- subtract 12 from each side 9x-9x=14-12 0=2 which is not right no answers

OpenStudy (anonymous):

thanks @GretaKnows

OpenStudy (anonymous):

did I do the math right?

OpenStudy (anonymous):

i need to make sure because state testing is this month

OpenStudy (jdoe0001):

well yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Yeah. SInce 9x-9x is 0. x does not equal anything therefore there is no solution

OpenStudy (jdoe0001):

another way to see it is y=3x–4 9x–3y=14 y=3x-4 <---- slope is "3" 9x -3y - 14 -3y = -9x - 14 y = 3x +(14/3) <---- slope is "3" same slope, lines are parallel, no solution, they never touch